# TAViS: Text-bridged Audio-Visual Segmentation with Foundation Models

### v1

@@ 1 Introduction
Audio-visual joint perception ability is crucial for our understanding of the surrounding environment. Various audio-visual collaboration tasks have emerged in recent years, including audio-visual sound separation [8, 38], visual sound source localization [4, 14, 15, 35], and audio-visual video understanding [16, 19, 23]. Distinct from these tasks, audio-visual segmentation (AVS) [47, 46] aims to generate pixel-wise segmentation maps of sounding objects within a scene, which has garnered increasing attention from the research community.
The primary challenge in AVS lies in effectively aligning audio and visual modalities, which requires capturing complex relationships between audio-visual patterns across diverse scenarios. Previous approaches are constrained by limited training data, resulting in incomplete capture of modal relationships. To mitigate this limitation, researchers have introduced large-scale datasets such as AVS-Synthetic [27] and VPO [7]. While pre-training on these datasets has shown promising improvements, it significantly increases computational overhead during training. Thanks to the generalization capabilities of foundation models, several works [44, 33, 27, 25] have utilized image foundation models, such as semantic-SAM [20] or SAM [6], as well as audio foundation models (e.g., Beats [5]), to leverage pre-existing knowledge with minimal additional training on AVS datasets. However, these approaches typically rely on single-modality foundation knowledge, either from audio or image, which hinders effective audio-visual modality alignment. While some previous works [3, 2] attempted to combine segmentation and multimodal foundation models, they employed these models in an off-the-shelf manner for unsupervised settings, where the models operated independently without meaningful interaction, as shown in Figure 1 (a). This reliance on a decoupled combination strategy compromises the transfer of foundation model knowledge.
In this paper, we propose a new model TAViS as shown in Figure 1 (b), designed to couple the knowledge provided by foundation models as a straightforward yet effective solution for the AVS task. However, choosing foundation models for the AVS task is challenging, as it requires both accurate recognition of sound-making objects and detailed segmentation of their boundaries. Our analysis leads us to focus on two complementary foundation models: ImageBind [10] for its robust audio-visual alignment, and SAM2 [36] for its strong generalization capability in complex image and video segmentation. To adapt SAM2 for object-level segmentation from mixed audio guidance, we design an ImageBind-guided query decomposition block (IBQD) that effectively separates audio sources while maintaining ImageBind’s well-aligned audio feature space.
@@ 2.1 Audio-Visual Segmentation
In recent years, AVS [45] task has garnered significant attention, as it addresses the challenging problem of locating and predicting pixel-wise maps for sounding objects within a scene. Existing AVS methods can be broadly categorized into three approaches: fusion-based, generative-based, and foundation model-based methods. Fusion-based methods constructed relationships between audio and visual inputs using unidirectional or bidirectional cross-attention mechanisms [45, 21, 9, 22, 44, 24]. For generative-based methods, latent diffusion models [29] or variational auto-encoders [30] were utilized to achieve accurate object segmentation. Foundation model-based methods leveraged visual foundation models such as SAM [18, 33, 27, 34, 37], UniDiffuser [25] and SAM-semantic [20] to obtain enhanced visual knowledge, facilitating the establishment of audio-visual relationships. However, these aforementioned foundation model-based methods primarily focus on visual prior knowledge, leaving a domain gap between audio and visual modalities unaddressed.
In this paper, we propose a novel approach that combines the strengths of Imagebind [10] and SAM2 [36]. Unlike BAVS [25], which utilizes text to align prediction maps, our method designs two separate losses to align text with visual and audio respectively, fully leveraging the text modality for cross-modal alignment. While [3, 2] also combined SAM and ImageBind for unsupervised audio-visual segmentation, they merely integrated these models in an off-the-shelf manner, which means the two models operate independently without interaction. In contrast, we introduce a text-bridged hybrid prompting design to align ImageBind with SAM2, thereby creating an organically coupled system that leads to more effective optimization.
@@ 2.2 Multimodal Joint Representation for Audio
Multimodal joint representation primarily focuses on mapping inputs from different modalities into a shared embedding space. In this section, we specifically focus on audio-related methods. For dual-modality alignment, WavCaps [31] pioneered multimodal representation learning for audio-text pairs, utilizing a large-scale weakly-labeled audio captioning dataset. In the multi-modal context, AudioCLIP [28] and WAV2CLIP [42] extended CLIP’s architecture by introducing dedicated audio encoders trained through contrastive learning, leveraging audio-image-text and audio-image pairs respectively. More recent approaches, including ImageBind [10], LanguageBind [48], and OmniBind [41], proposed comprehensive frameworks for integrating multiple modalities by utilizing various data pairs that share common image, language, and text modalities.
Existing AVS methods [44, 33, 27, 25] typically rely on single-modality foundation models for either audio or visual processing independently, resulting in heterogeneous feature spaces that complicate cross-modal alignment. While [32] trained a specialized foundation model for audio-visual tasks, it lacks text alignment capabilities and predicts fixed classes via MLP prediction, which constrains its applicability in real-world scenarios. In contrast to these approaches, our model can perform zero-shot setting, demonstrating superior generalizability for unseen classes through our innovative text-bridged design approach.
@@ 3 Methodology
In this work, we propose TAViS, a novel framework that leverages foundation models to achieve precise region segmentation by bridging audio and visual content through text representations. Built upon SAM2 [36] and ImageBind [12] (Section 3.1), our approach introduces three key components: 1) an ImageBind-guided query decomposition module to decompose the mixed audio embedding obtained from ImageBind into object-level queries (Section 3.2).
@@ 3.1 Preliminaries
For visual segmentation, we adopt SAM2 [36] as our foundation model. SAM2 consists of three key components: an image encoder $F_{E}$ to extract visual features, a memory attention module $F_{M}$ to maintain temporal consistency, and a mask decoder $F_{D}$ to generate precise segmentation masks for regions of interest. To establish effective audio-visual alignment, we incorporate ImageBind [12] into our framework. Within the ImageBind architecture, we specifically utilize its image encoder $E_{I}$, text encoder $E_{T}$, and audio encoder $E_{A}$. For each modality encoder, we obtain the trunk feature and a cls token with a transformer architecture.
Given an audio signal $\bm{A}$ and its corresponding video frames $\bm{I}\in{\mathbb{R}^{T\times H\times W\times C}}$, where $T$ denotes the total number of frames in the video, we first extract visual features through SAM2’s image encoder $F_{E}$, generating multi-stage visual features. Concurrently, we process the audio input $\bm{A}$ through ImageBind’s audio encoder $E_{A}$ to obtain the audio trunk feature $\bm{f}_{a}$ and the cls token $\bm{t}_{a}$. To enhance the correlation between audio and visual modalities, we introduce audio-guided adaptors at each stage of SAM2’s visual encoder $F_{E}$, following the approach in [27].
@@ 3.2 ImageBind-guided Query Decomposition
Previous SAM-based AVS methods [18, 33, 27, 34] typically treat the entire audio feature as a single prompt for the mask decoder. However, the mixing of information from multiple sound sources conflicts with the object-specific queries required by SAM2. A straightforward way to mitigate this issue is to decompose the entire audio feature to object-level prompts with a learnable decomposition module. However, naively doing this can not leverage the well-aligned multimodal knowledge learned by ImageBind. To address this challenge, we propose an ImageBind-guided Query Decomposition (IBQD) to decompose audio features into object-specific queries that preserve the aligned feature space of ImageBind, enabling more precise subsequent audio-visual correspondence.
@@
For simplicity, we omit the layer normalization [1] operation here. Subsequently, a linear transformation is applied to $\bm{t}_{W}$ to generate object-specific biases, which are added to $\bm{t}_{a}$ and obtain $\bm{t}^{\prime}_{a}\in{\mathbb{R}^{N\times C}}$. Finally, the $\bm{t}^{\prime}_{a}$ is updated by $\bm{f}_{a}$ again with MCA:
@@ 3.3 Text-bridged Hybrid Prompting
Prompts play a crucial role in SAM2, providing essential guidance to specify the object of interest to segment. Previous works based on SAM [18, 33, 27, 34] typically process audio inputs through learnable MLPs to generate semantic prompts. However, they may struggle to capture the inherent class prototype nature of audio information required for effective AVS alignment. In AVS scenarios, the initial audio input contains substantial non-semantic information (such as diverse timbres and tones) that can confuse the model. Therefore, we propose novel text-bridged hybrid prompts including sparse prompt and dense prompt, where text provides class prototype information while audio and visual modalities complement with modality-specific details. The overall architecture is shown in Figure 3.
@@
For sparse prompt, we adopt an audio-text dual prompt that leverages text and audio to provide the general object concepts and specific instance details, respectively. We first generate specific object-level text prototype information and scene-level global context following [43]. Specifically, we utilize learnable MLP to the decomposed audio queries $\bm{t}^{\prime}_{a}$ and the audio cls token $\bm{t}_{a}$. These two transformed tokens are then concatenated and processed by the text encoder $E_{T}$ to generate the pseudo text prompt $\bm{p}^{t}$.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our model on two AVSBench datasets: AVSBench-object [47] and AVSBench-semantic [46]. The AVSBench-object dataset comprises two scenarios: single sound source segmentation (S4) with 4,932 videos and multiple sound source segmentation (MS3) with 424 videos. MS3 contains 5 annotated frames per video, while S4 requires semi-supervised training, as only the first frame is annotated out of 5 frames. The AVSBench-semantic dataset introduces two key differences compared with AVSBench-object. First, it adds a new scenario (V2) containing 6,000 videos with 10 frames per audio clip. Second, it enhances the S4 and MS3 subsets with object-level class label, creating new datasets V1S and V1M, respectively.
@@
We adopt two segmentation metrics: $\mathcal{M}_{\mathcal{J}}$ which measures the overlap between ground-truth masks and predicted ones, and $\mathcal{M}_{\mathcal{F}}$ which considers both precision and recall in the evaluation. Following previous work [44, 21, 47, 26], we use $\mathcal{M}_{\mathcal{J}}$ and $\mathcal{M}_{\mathcal{F}}$ to denote the mean metrics across the entire dataset. On S4 and MS3, we evaluate binary segmentation performance while on AVSS, we report results based on semantic-level metric $\mathcal{M}^{I}_{\mathcal{J}}$.
@@
$\mathcal{L}_{a2t}$ +$\mathcal{L}_{i2t}$ + $\mathcal{L}_{a2i}$ 83.6 0.903 66.5 0.744
@@
      AVSBench [45]       PVT-v2       224       78.7       0.879       54.0       0.645       -
      CATR [21]       PVT-v2       224       81.4       0.896       59.0       0.700       -
      DiffusionAVS [29]       PVT-v2       224       81.4       0.902       58.2       0.709       -
      ECMVAE [30]       PVT-v2       224       81.7       0.901       57.8       0.708       -
      BAVS [25]       PVT-v2       224       82.0       0.886       58.6       0.655       32.6
      AVSegFormer [9]       PVT-v2       224       82.1       0.899       58.4       0.693       36.7
      UFE-AVS [26]       PVT-v2       224       83.2       0.904       59.5       0.676       -
      QDFormer [22]       Swin-T       224       79.5       0.882       61.9       0.661       -
      COMBO [44]       PVT-v2       224       84.7       0.919       59.2       0.712       42.1
      SAM [18]       ViT-H       1024       55.1       0.739       54.0       0.638       -
      AV-SAM [33]       ViT-H       1024       40.5       0.566       -       -       -
      GAVS [40]       ViT-B       1024       80.1       0.902       63.7       0.774       -
      SAMA-AVS [27]       ViT-H       1024       83.2       0.901       66.9       0.754       -
@@ 4.2 Implementation Details
We utilize the ViT-L backbone for SAM2 and freeze the parameters of ImageBind and the SAM2 image encoder, while fine-tuning the memory attention and mask decoder. All input frames are resized to 224×224 or 1024×1024. Training is performed using the Adam optimizer [17], starting with an initial learning rate of 1e-4 and a cosine decay scheduler. The batch sizes used are 4 for the MS3 and AVSS datasets, and 10 for the S4 dataset. Training lasts for 80 epochs on both the MS3 and S4 datasets, and 40 epochs on the AVSS dataset.

### v2

@@ 1 Introduction
Audio-visual joint perception ability is crucial for our understanding of the surrounding environment. Various audio-visual collaboration tasks have emerged in recent years, including audio-visual sound separation [gao2021visualvoice, tzinis2022audioscopev2], visual sound source localization [chen2021localizing, hu2020discriminative, hu2021class, qian2020multiple], and audio-visual video understanding [kazakos2019epic, lee2020cross, lin2019dual]. Distinct from these tasks, audio-visual segmentation (AVS) [zhou2022audio, zhou2024audio] aims to generate pixel-wise segmentation maps of sounding objects within a scene, which has garnered increasing attention from the research community.
The primary challenge in AVS lies in effectively aligning audio and visual modalities, which requires capturing complex relationships between audio-visual patterns across diverse scenarios. Previous approaches are constrained by limited training data, resulting in incomplete capture of modal relationships. To mitigate this limitation, researchers have introduced large-scale datasets such as AVS-Synthetic [liu2024annotation] and VPO [chen2024unraveling]. While pre-training on these datasets has shown promising improvements, it significantly increases computational overhead during training. Thanks to the generalization capabilities of foundation models, several works [yang2024cooperation, mo2023av, liu2024annotation, liu2024bavs] have utilized image foundation models, such as semantic-SAM [li2023semantic] or SAM [chen2023sam], as well as audio foundation models (e.g., Beats [chen2022beats]), to leverage pre-existing knowledge with minimal additional training on AVS datasets. However, these approaches typically rely on single-modality foundation knowledge, either from audio or image, which hinders effective audio-visual modality alignment. While some previous works [bhosale2023leveraging, bhosale2024unsupervised] attempted to combine segmentation and multimodal foundation models, they employed these models in an off-the-shelf manner for unsupervised settings, where the models operated independently without meaningful interaction, as shown in Figure 1 (a). This reliance on a decoupled combination strategy compromises the transfer of foundation model knowledge.
In this paper, we propose a new model TAViS as shown in Figure 1 (b), designed to couple the knowledge provided by foundation models as a straightforward yet effective solution for the AVS task. However, choosing foundation models for the AVS task is challenging, as it requires both accurate recognition of sound-making objects and detailed segmentation of their boundaries. Our analysis leads us to focus on two complementary foundation models: ImageBind [girdhar2023imagebind] for its robust audio-visual alignment, and SAM2 [ravi2024sam] for its strong generalization capability in complex image and video segmentation. To adapt SAM2 for object-level segmentation from mixed audio guidance, we design an ImageBind-guided query decomposition block (IBQD) that effectively separates audio sources while maintaining ImageBind’s well-aligned audio feature space.
@@ 2.1 Audio-Visual Segmentation
In recent years, AVS [zhou2023audio] task has garnered significant attention, as it addresses the challenging problem of locating and predicting pixel-wise maps for sounding objects within a scene. Existing AVS methods can be broadly categorized into three approaches: fusion-based, generative-based, and foundation model-based methods. Fusion-based methods constructed relationships between audio and visual inputs using unidirectional or bidirectional cross-attention mechanisms [zhou2023audio, li2023catr, gao2024avsegformer, li2024qdformer, yang2024cooperation, ling2024transavs]. For generative-based methods, latent diffusion models [mao2023contrastive] or variational auto-encoders [mao2023multimodal] were utilized to achieve accurate object segmentation. Foundation model-based methods leveraged visual foundation models such as SAM [kirillov2023segment, mo2023av, liu2024annotation, nguyen2024save, seon2024extending], UniDiffuser [liu2024bavs] and SAM-semantic [li2023semantic] to obtain enhanced visual knowledge, facilitating the establishment of audio-visual relationships. However, these aforementioned foundation model-based methods primarily focus on visual prior knowledge, leaving a domain gap between audio and visual modalities unaddressed.
In this paper, we propose a novel approach that combines the strengths of Imagebind [girdhar2023imagebind] and SAM2 [ravi2024sam]. Unlike BAVS [liu2024bavs], which utilizes text to align prediction maps, our method designs two separate losses to align text with visual and audio respectively, fully leveraging the text modality for cross-modal alignment. While [bhosale2023leveraging, bhosale2024unsupervised] also combined SAM and ImageBind for unsupervised audio-visual segmentation, they merely integrated these models in an off-the-shelf manner, which means the two models operate independently without interaction. In contrast, we introduce a text-bridged hybrid prompting design to align ImageBind with SAM2, thereby creating an organically coupled system that leads to more effective optimization.
@@ 2.2 Multimodal Joint Representation for Audio
Multimodal joint representation primarily focuses on mapping inputs from different modalities into a shared embedding space. In this section, we specifically focus on audio-related methods. For dual-modality alignment, WavCaps [mei2024wavcaps] pioneered multimodal representation learning for audio-text pairs, utilizing a large-scale weakly-labeled audio captioning dataset. In the multi-modal context, AudioCLIP [mahmud2023ave] and WAV2CLIP [wu2022wav2clip] extended CLIP’s architecture by introducing dedicated audio encoders trained through contrastive learning, leveraging audio-image-text and audio-image pairs respectively. More recent approaches, including ImageBind [girdhar2023imagebind], LanguageBind [zhu2023languagebind], and OmniBind [wang2024omnibind], proposed comprehensive frameworks for integrating multiple modalities by utilizing various data pairs that share common image, language, and text modalities.
Existing AVS methods [yang2024cooperation, mo2023av, liu2024annotation, liu2024bavs] typically rely on single-modality foundation models for either audio or visual processing independently, resulting in heterogeneous feature spaces that complicate cross-modal alignment. While [mo2024unveiling] trained a specialized foundation model for audio-visual tasks, it lacks text alignment capabilities and predicts fixed classes via MLP prediction, which constrains its applicability in real-world scenarios. In contrast to these approaches, our model can perform zero-shot setting, demonstrating superior generalizability for unseen classes through our innovative text-bridged design approach.
@@ 3 Methodology
In this work, we propose TAViS, a novel framework that leverages foundation models to achieve precise region segmentation by bridging audio and visual content through text representations. Built upon SAM2 [ravi2024sam] and ImageBind [han2023imagebind] (Section 3.1), our approach introduces three key components: 1) an ImageBind-guided query decomposition module to decompose the mixed audio embedding obtained from ImageBind into object-level queries (Section 3.2).
@@ 3.1 Preliminaries
For visual segmentation, we adopt SAM2 [ravi2024sam] as our foundation model. SAM2 consists of three key components: an image encoder $F_{E}$ to extract visual features, a memory attention module $F_{M}$ to maintain temporal consistency, and a mask decoder $F_{D}$ to generate precise segmentation masks for regions of interest. To establish effective audio-visual alignment, we incorporate ImageBind [han2023imagebind] into our framework. Within the ImageBind architecture, we specifically utilize its image encoder $E_{I}$, text encoder $E_{T}$, and audio encoder $E_{A}$. For each modality encoder, we obtain the trunk feature and a cls token with a transformer architecture.
Given an audio signal $\bm{A}$ and its corresponding video frames $\bm{I}\in{\mathbb{R}^{T\times H\times W\times C}}$, where $T$ denotes the total number of frames in the video, we first extract visual features through SAM2’s image encoder $F_{E}$, generating multi-stage visual features. Concurrently, we process the audio input $\bm{A}$ through ImageBind’s audio encoder $E_{A}$ to obtain the audio trunk feature $\bm{f}_{a}$ and the cls token $\bm{t}_{a}$. To enhance the correlation between audio and visual modalities, we introduce audio-guided adaptors at each stage of SAM2’s visual encoder $F_{E}$, following the approach in [liu2024annotation].
@@ 3.2 ImageBind-guided Query Decomposition
Previous SAM-based AVS methods [kirillov2023segment, mo2023av, liu2024annotation, nguyen2024save] typically treat the entire audio feature as a single prompt for the mask decoder. However, the mixing of information from multiple sound sources conflicts with the object-specific queries required by SAM2. A straightforward way to mitigate this issue is to decompose the entire audio feature to object-level prompts with a learnable decomposition module. However, naively doing this can not leverage the well-aligned multimodal knowledge learned by ImageBind. To address this challenge, we propose an ImageBind-guided Query Decomposition (IBQD) to decompose audio features into object-specific queries that preserve the aligned feature space of ImageBind, enabling more precise subsequent audio-visual correspondence.
@@
For simplicity, we omit the layer normalization [ba2016layer] operation here. Subsequently, a linear transformation is applied to $\bm{t}_{W}$ to generate object-specific biases, which are added to $\bm{t}_{a}$ and obtain $\bm{t}^{\prime}_{a}\in{\mathbb{R}^{N\times C}}$. Finally, the $\bm{t}^{\prime}_{a}$ is updated by $\bm{f}_{a}$ again with MCA:
@@ 3.3 Text-bridged Hybrid Prompting
Prompts play a crucial role in SAM2, providing essential guidance to specify the object of interest to segment. Previous works based on SAM [kirillov2023segment, mo2023av, liu2024annotation, nguyen2024save] typically process audio inputs through learnable MLPs to generate semantic prompts. However, they may struggle to capture the inherent class prototype nature of audio information required for effective AVS alignment. In AVS scenarios, the initial audio input contains substantial non-semantic information (such as diverse timbres and tones) that can confuse the model. Therefore, we propose novel text-bridged hybrid prompts including sparse prompt and dense prompt, where text provides class prototype information while audio and visual modalities complement with modality-specific details. The overall architecture is shown in Figure 3.
@@
For sparse prompt, we adopt an audio-text dual prompt that leverages text and audio to provide the general object concepts and specific instance details, respectively. We first generate specific object-level text prototype information and scene-level global context following [wu2023aligning]. Specifically, we utilize learnable MLP to the decomposed audio queries $\bm{t}^{\prime}_{a}$ and the audio cls token $\bm{t}_{a}$. These two transformed tokens are then concatenated and processed by the text encoder $E_{T}$ to generate the pseudo text prompt $\bm{p}^{t}$.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our model on two AVSBench datasets: AVSBench-object [zhou2022audio] and AVSBench-semantic [zhou2024audio]. The AVSBench-object dataset comprises two scenarios: single sound source segmentation (S4) with 4,932 videos and multiple sound source segmentation (MS3) with 424 videos. MS3 contains 5 annotated frames per video, while S4 requires semi-supervised training, as only the first frame is annotated out of 5 frames. The AVSBench-semantic dataset introduces two key differences compared with AVSBench-object. First, it adds a new scenario (V2) containing 6,000 videos with 10 frames per audio clip. Second, it enhances the S4 and MS3 subsets with object-level class label, creating new datasets V1S and V1M, respectively.
@@
We adopt two segmentation metrics: $\mathcal{M}_{\mathcal{J}}$ which measures the overlap between ground-truth masks and predicted ones, and $\mathcal{M}_{\mathcal{F}}$ which considers both precision and recall in the evaluation. Following previous work [yang2024cooperation, li2023catr, zhou2022audio, liu2024audio], we use $\mathcal{M}_{\mathcal{J}}$ and $\mathcal{M}_{\mathcal{F}}$ to denote the mean metrics across the entire dataset. On S4 and MS3, we evaluate binary segmentation performance while on AVSS, we report results based on semantic-level metric $\mathcal{M}^{I}_{\mathcal{J}}$.
@@
$\mathcal{L}_{a2t}$ $\mathcal{L}_{i2t}$ + $\mathcal{L}_{a2i}$ 83.6 0.903 66.5 0.744
@@
   AVSBench [zhou2023audio]    PVT-v2    224    78.7    0.879    54.0    0.645    -
   CATR [li2023catr]    PVT-v2    224    81.4    0.896    59.0    0.700    -
   DiffusionAVS [mao2023contrastive]    PVT-v2    224    81.4    0.902    58.2    0.709    -
   ECMVAE [mao2023multimodal]    PVT-v2    224    81.7    0.901    57.8    0.708    -
   BAVS [liu2024bavs]    PVT-v2    224    82.0    0.886    58.6    0.655    32.6
   AVSegFormer [gao2024avsegformer]    PVT-v2    224    82.1    0.899    58.4    0.693    36.7
   UFE-AVS [liu2024audio]    PVT-v2    224    83.2    0.904    59.5    0.676    -
   QDFormer [li2024qdformer]    Swin-T    224    79.5    0.882    61.9    0.661    -
   COMBO [yang2024cooperation]    PVT-v2    224    84.7    0.919    59.2    0.712    42.1
   SAM [kirillov2023segment]    ViT-H    1024    55.1    0.739    54.0    0.638    -
   AV-SAM [mo2023av]    ViT-H    1024    40.5    0.566    -    -    -
   GAVS [wang2024prompting]    ViT-B    1024    80.1    0.902    63.7    0.774    -
   SAMA-AVS [liu2024annotation]    ViT-H    1024    83.2    0.901    66.9    0.754    -
@@ 4.2 Implementation Details
We utilize the ViT-L backbone for SAM2 and freeze the parameters of ImageBind and the SAM2 image encoder, while fine-tuning the memory attention and mask decoder. All input frames are resized to 224×224 or 1024×1024. Training is performed using the Adam optimizer [kingma2014adam], starting with an initial learning rate of 1e-4 and a cosine decay scheduler. The batch sizes used are 4 for the MS3 and AVSS datasets, and 10 for the S4 dataset. Training lasts for 80 epochs on both the MS3 and S4 datasets, and 40 epochs on the AVSS dataset.